# MeanFuser: Fast One-Step Multi-Modal Trajectory Generation and Adaptive Reconstruction via MeanFlow for End-to-End Autonomous Driving

### v1

@@ Abstract
Generative models have shown great potential in trajectory planning. Recent studies demonstrate that anchor-guided generative models are effective in modeling the uncertainty of driving behaviors and improving overall performance. However, these methods rely on discrete anchor vocabularies that must sufficiently cover the trajectory distribution during testing to ensure robustness, inducing an inherent trade-off between vocabulary size and model performance. To overcome this limitation, we propose MeanFuser, an end-to-end autonomous driving method that enhances both efficiency and robustness through three key designs. (1) We introduce Gaussian Mixture Noise (GMN) to guide generative sampling, enabling a continuous representation of the trajectory space and eliminating the dependency on discrete anchor vocabularies. (2) We adapt “MeanFlow Identity” to end-to-end planning, which models the mean velocity field between GMN and trajectory distribution instead of the instantaneous velocity field used in vanilla flow matching methods, effectively eliminating numerical errors from ODE solvers and significantly accelerating inference. (3) We design a lightweight Adaptive Reconstruction Module (ARM) that enables the model to implicitly select from all sampled proposals or reconstruct a new trajectory when none is satisfactory via attention weights. Experiments on the NAVSIM closed-loop benchmark demonstrate that MeanFuser achieves outstanding performance without the supervision of the PDM Score. and exceptional inference efficiency, offering a robust and efficient solution for end-to-end autonomous driving. Our code and model are available at https://github.com/wjl2244/MeanFuser.
footnotetext: *Completed during internship at Xiaomi EV. † Corresponding author
@@ 1 Introduction
End-to-end autonomous driving[2, 30, 5, 11, 12, 13, 19, 22, 36] has attracted considerable attention as it directly learns the mapping from raw sensor inputs to planned trajectories, thereby avoiding handcrafted rules in rule-based systems and the cumulative errors inherent in modular pipelines.
@@ 2.1 End-to-End Autonomous Driving
End-to-end autonomous driving systems take sensory inputs and ego-vehicle states as inputs and aim to directly generate driving trajectories through neural optimization. Early works such as Transfuser[5] focused on fusing heterogeneous modalities such as camera and LiDAR, while introducing auxiliary tasks to enhance supervision during action-space training. UniAD[12] further structured the entire framework as a planning-oriented transformer pipeline, enabling information flow across multiple specialized modules. Building on this line of research, VAD[13] replaced dense bird’s-eye-view representations with lightweight vectorized scene abstractions, which significantly improved efficiency. Its successor, VADv2[3], reformulated trajectory regression as a discrete trajectory-token selection problem, choosing the optimal trajectory from a learned vocabulary. ParaDrive[29] departed from the traditional serial end-to-end paradigm and adopted a parallel framework, allowing mapping, prediction, and planning to proceed simultaneously. Unlike the aforementioned methods, we use powerful generative models to model the trajectory distribution.
@@ 2.2 Diffusion and Flow-Based Generative Models for Trajectory Planning
Driving behaviors are inherently multimodal since in complex traffic scenarios, multiple feasible trajectories may exist. To capture this stochasticity, recent works adopt diffusion[10, 24, 25] or flow matching[20, 21] generative paradigms, training with noise injection and denoising inference to sample diverse yet plausible trajectories[19, 27, 31]. MotionDiffuser[14] was among the first to introduce diffusion-based trajectory generation in autonomous driving. Later, Diffusion-ES[32] refined the generated trajectories through iterative optimization that alternates between generation and evaluation. Diffusion Planner[37] incorporated scene context through carefully designed conditioning mechanisms that effectively guided the diffusion process. Moreover, DiffusionDrive[19] introduced noise perturbations to clustered trajectory prototypes, effectively encoding prior trajectory distributions. In contrast, GoalFlow[31] employed flow matching for improved stability and utilized goal points to explicitly guide trajectory generation. We improve the model’s performance and robustness without relying on anchor-based guidance by introducing Gaussian Mixture Noise sampling and an Adaptive Reconstruction Module, while requiring only one-step sampling.
@@ 2.3 Candidate Trajectory Proposal Evaluation and Selection
Figure 3: Overall architecture of MeanFuser. Training: During training, both the images and ego-vehicle states are encoded into context features, with auxiliary supervision from mapping and detection tasks. The model is conditioned on these context features to learn the average velocity field $u_{\theta}$ over the time interval $r$ and $t$. Multi-Modal Sample: Noise samples are drawn from Gaussian Mixture Noise, and the one-step sampling formulation is then applied to generate diverse multi-modal trajectories. Adaptive Reconstruction Module: The sampled multi-modal trajectories are encoded and fused with the context features through cross-attention, after which a Projector outputs the final planning trajectory.
@@ 4.1 Scene Context Encoder
During training, an auxiliary decoder is employed to decode dynamic information of surrounding traffic participants and the lane map from $\mathbf{c_{bev}}$, with the corresponding loss functions $\mathcal{L}_{\text{det}}$ and $\mathcal{L}_{\text{map}}$ applied to supervise semantic learning and accelerate model convergence. The extracted scene features $\mathbf{c}=\{\mathbf{c_{bev}},\mathbf{c_{s}}\}$ serve as scene-specific conditional inputs for multi-modal trajectories sampling.
@@ 4.2 Gaussian Mixture Noise
UniAD[12] CVPR 2023 C & L ResNet-34 97.8 91.9 92.9 100 78.8 83.4
PARADrive[29] CVPR 2024 C & L ResNet-34 97.9 92.4 93.0 99.8 79.3 84.0
@@
MeanFuser(Ours) - C-Only ResNet-34 98.6 97.0 95.0 100 82.8 89.0
@@ 4.4 Adaptive Reconstruction Module
The total loss function is a weighted sum of the losses from Meanflow decoder, Mapping, Detection, and ARM, as follows:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\tau}+\lambda_{2}\mathcal{L}_{flow}+\lambda_{3}\mathcal{L}_{map}+\lambda_{4}\mathcal{L}_{det}$ (15)
The hyperparameters $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ and $\lambda_{4}$ control the relative importance of each component loss.

### v2

@@ Abstract
Generative models have shown great potential in trajectory planning. Recent studies demonstrate that anchor-guided generative models are effective in modeling the uncertainty of driving behaviors and improving overall performance. However, these methods rely on discrete anchor vocabularies that must sufficiently cover the trajectory distribution during testing to ensure robustness, inducing an inherent trade-off between vocabulary size and model performance. To overcome this limitation, we propose MeanFuser, an end-to-end autonomous driving method that enhances both efficiency and robustness through three key designs. (1) We introduce Gaussian Mixture Noise (GMN) to guide generative sampling, enabling a continuous representation of the trajectory space and eliminating the dependency on discrete anchor vocabularies. (2) We adapt “MeanFlow Identity” to end-to-end planning, which models the mean velocity field between GMN and trajectory distribution instead of the instantaneous velocity field used in vanilla flow matching methods, effectively eliminating numerical errors from ODE solvers and significantly accelerating inference. (3) We design a lightweight Adaptive Reconstruction Module (ARM) that enables the model to implicitly select from all sampled proposals or reconstruct a new trajectory when none is satisfactory via attention weights.Experiments on the NAVSIM closed-loop benchmark demonstrate that MeanFuser achieves outstanding performance without the supervision of the PDM Score and exceptional inference efficiency, offering a robust and efficient solution for end-to-end autonomous driving.
footnotetext: † Corresponding author. This work is supported by the Beijing Natural Science Foundation-Xiaomi Innovation Joint Fund L253007, and Beijing Natural Science Foundation under Grant 4242052. Our code are available at https://github.com/wjl2244/MeanFuser.
@@ 1 Introduction
End-to-end autonomous driving[2, 7, 13, 25, 39] has attracted considerable attention as it directly learns the mapping from raw sensor inputs to planned trajectories, thereby avoiding handcrafted rules in rule-based systems and the cumulative errors inherent in modular pipelines[41].
@@ 2.1 End-to-End Autonomous Driving
End-to-end autonomous driving systems take sensory inputs and ego-vehicle states as inputs and aim to directly generate driving trajectories through neural optimization[42, 35]. Early works such as Transfuser[7] focused on fusing heterogeneous modalities such as camera and LiDAR, while introducing auxiliary tasks to enhance supervision during action-space training. UniAD[14] further structured the entire framework as a planning-oriented transformer pipeline, enabling information flow across multiple specialized modules. Building on this line of research, VAD[15] replaced dense bird’s-eye-view representations with lightweight vectorized scene abstractions, which significantly improved efficiency. Its successor, VADv2[3], reformulated trajectory regression as a discrete trajectory-token selection problem, choosing the optimal trajectory from a learned vocabulary. ParaDrive[32] departed from the traditional serial end-to-end paradigm and adopted a parallel framework, allowing mapping, prediction, and planning to proceed simultaneously. Unlike the aforementioned methods, we use powerful generative models to model the trajectory distribution.
@@ 2.2 Diffusion and Flow-Based Generative Models for Trajectory Planning
Driving behaviors are inherently multimodal since in complex traffic scenarios, multiple feasible trajectories may exist. To capture this stochasticity, recent works adopt diffusion[4, 17, 5] or flow matching[23, 24] generative paradigms, training with noise injection and denoising inference to sample diverse yet plausible trajectories[22, 30, 10, 33]. MotionDiffuser[16] was among the first to introduce diffusion-based trajectory generation in autonomous driving. Later, Diffusion-ES[34] refined the generated trajectories through iterative optimization that alternates between generation and evaluation. Diffusion Planner[40] incorporated scene context through carefully designed conditioning mechanisms that effectively guided the diffusion process. Moreover, DiffusionDrive[22] introduced noise perturbations to clustered trajectory prototypes, effectively encoding prior trajectory distributions. In contrast, GoalFlow[33] employed flow matching for improved stability and utilized goal points to explicitly guide trajectory generation. We improve the model’s performance and robustness without relying on anchor-based guidance by introducing Gaussian Mixture Noise sampling and an Adaptive Reconstruction Module, while requiring only one-step sampling.
@@ 2.3 Candidate Trajectory Proposal Evaluation and Selection
Figure 3: Overall architecture of MeanFuser. Training: During training, both the images and ego-vehicle states are encoded into context features, with auxiliary supervision from mapping tasks. The model is conditioned on these context features to learn the average velocity field $u_{\theta}$ over the time interval $r$ and $t$. Multi-Modal Sample: Noise samples are drawn from Gaussian Mixture Noise, and the one-step sampling formulation is then applied to generate diverse multi-modal trajectories. Adaptive Reconstruction Module: The sampled multi-modal trajectories are encoded and fused with the context features through cross-attention, after which a Projector outputs the final planning trajectory.
@@ 4.1 Scene Context Encoder
During training, an auxiliary decoder is employed to decode information of the lane map from $\mathbf{c_{bev}}$, with the corresponding loss functions $\mathcal{L}_{\text{map}}$ applied to supervise semantic learning and accelerate model convergence. The extracted scene features $\mathbf{c}=\{\mathbf{c_{bev}},\mathbf{c_{s}}\}$ serve as scene-specific conditional inputs for multi-modal trajectories sampling.
@@ 4.2 Gaussian Mixture Noise
UniAD[14] CVPR 2023 C ResNet-34 97.8 91.9 92.9 100 78.8 83.4
World4Drive[42] ICCV 2025 C ResNet-34 97.4 94.3 92.8 100 79.9 85.1
Epona [37] ICCV 2025 C Transformer 97.9 95.1 93.8 99.9 80.4 86.2
MeanFuser(Ours) - C ResNet-34 98.6 97.0 95.0 100 82.8 89.0
@@ 4.4 Adaptive Reconstruction Module
The total loss function is a weighted sum of the losses from Meanflow decoder, Mapping and ARM, as follows:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\tau}+\lambda_{2}\mathcal{L}_{flow}+\lambda_{3}\mathcal{L}_{map}$ (15)
The hyperparameters $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ control the relative importance of each component loss.